# gRNAde: Geometric Deep Learning for 3D RNA inverse design

### v4

@@ Appendix C Additional Figures and Tables
Table 1: Full results for Figure 4 comparing gRNAde to Rosetta and FARNA for single-state design.
Figure 9: RNASolo data statistics.
Table 2: Aggregated benchmark results for gRNAde.
@@
Figure 8: In-silico evaluation metrics for gRNAde designed sequences. We consider (1) the percentage of native nucleotides recovered in designed samples, (2) self-consistency of ‘forward folded’ secondary structures of designs with the native secondary structure, as well as (3) perplexity of the model’s prediction (not shown in the figure).
Table 1: Full results for Figure 4 comparing gRNAde to Rosetta and FARNA for single-state design on 14 RNA structures of interest identified by Das et al. (2010). Rosetta and FARNA recovery values are taken from Das et al. (2010), Supplementary Table 2.
FARNA Rosetta gRNAde (single-state)
PDB ID Description Recovery Recovery Recovery Perplexity Self-cons.
1CSL RRE high affinity site 0.20 0.44 0.5719 1.2812 0.8644
1ET4 Vitamin B12 binding RNA aptamer 0.34 0.44 0.6250 1.3457 -0.0135
1F27 Biotin-binding RNA pseudoknot 0.36 0.37 0.3437 1.6203 0.4523
1L2X Viral RNA pseudoknot 0.45 0.48 0.4721 1.3181 0.5692
1LNT RNA internal loop of SRP 0.27 0.53 0.5843 1.4337 0.1379
1Q9A Sarcin/ricin domain from E.coli 23S rRNA 0.40 0.41 0.5044 1.3411 0.0597
4FE5 Guanine riboswitch aptamer 0.28 0.36 0.5300 1.3824 0.9116
1X9C All-RNA hairpin ribozyme 0.31 0.50 0.5000 1.3905 0.6630
1XPE HIV-1 B RNA dimerization initiation site 0.24 0.40 0.7037 1.2177 0.7768
2GCS Pre-cleavage state of glmS ribozyme 0.26 0.44 0.5078 1.3053 0.4062
2GDI Thiamine pyrophosphate-specific riboswitch 0.38 0.48 0.6500 1.2363 -0.0251
2OEU Junctionless hairpin ribozyme 0.30 0.37 0.9519 1.0913 0.7768
2R8S Synthetic FAB bound to ribozyme domain 0.36 0.53 0.5689 1.1881 0.7281
354D Loop E from E. coli 5S rRNA 0.35 0.55 0.4410 1.4938 0.0430
Overall recovery: 0.32 0.45 0.5682
Table 2: Aggregated benchmark results for gRNAde. Split: Single- and multi-state splits are described in Section 3; models trained on ‘All data’ use all RNASolo samples for training and are evaluated on the multi-state test set, solely for the purpose of releasing the best possible gRNAde checkpoints for real-world usage. Model: ‘AR’ implies autoregressive decoding (described in Section 2.3), while ‘NAR’ implies non-autoregressive, one-shot decoding using an MLP. AR models have significantly higher self-consistency scores than NAR, as autoregressive decoding can condition predictions at each iteration on past predictions while one-shot decoding cannot. Max. #states: Multi-state models marginally improve native sequence recovery over an equivalent single state variant, even for the single-state benchmark. Max. train RNA length: Limiting the maximum length of RNAs used for training often leads to lower sequence recovery and higher self-consistency scores. Results in the main paper are reported for models shaded in gray which are all trained on maximum length of 5000 nucleotides.
Max. Max. train Native seq. Sec. struct. MCC
Split Model #states RNA length recovery self-consistency
Single-state split AR 1 500 0.4364$\pm$0.0059 0.6206$\pm$0.0662
AR 1 1000 0.4534$\pm$0.0063 0.6481$\pm$0.0037
AR 1 2500 0.4945$\pm$0.0077 0.6278$\pm$0.0215
AR 1 5000 0.5271$\pm$0.0019 0.5706$\pm$0.0119
AR 3 5000 0.5386$\pm$0.0134 0.6255$\pm$0.0057
AR 5 5000 0.5400$\pm$0.0270 0.6006$\pm$0.0345
Multi-state split AR 1 500 0.4452$\pm$0.0067 0.6031$\pm$0.0288
AR 1 1000 0.4472$\pm$0.0132 0.5799$\pm$0.0156
AR 1 2000 0.4799$\pm$0.0180 0.5668$\pm$0.0120
AR 1 5000 0.4549$\pm$0.0082 0.5689$\pm$0.0185
AR 3 5000 0.4666$\pm$0.0288 0.5605$\pm$0.0305
AR 5 5000 0.4722$\pm$0.0106 0.5488$\pm$0.0323
NAR 1 5000 0.4324$\pm$0.0097 0.4229$\pm$0.0050
NAR 3 5000 0.4327$\pm$0.0212 0.4185$\pm$0.0249
NAR 5 5000 0.4450$\pm$0.0116 0.4130$\pm$0.0177
All data AR 1 5000 0.7361$\pm$0.0039 0.6301$\pm$0.0160
AR 2 5000 0.7845$\pm$0.0099 0.6215$\pm$0.0218
AR 3 5000 0.7886$\pm$0.0106 0.6052$\pm$0.0128
AR 5 5000 0.8110$\pm$0.0077 0.6135$\pm$0.0280

### v5

@@ Appendix E Additional Results
Table 2: Full results for Figure 4 comparing gRNAde to Rosetta, FARNA and ViennaRNA for single-state design on 14 RNA structures of interest identified by Das et al. (2010). Rosetta and FARNA recovery values are taken from Das et al. (2010), Supplementary Table 2.
ViennaRNA FARNA Rosetta gRNAde (single-state)
PDB ID Description Recovery Recovery Recovery Recovery Perplexity 2D self-cons.
1CSL RRE high affinity site 0.25 0.20 0.44 0.5719 1.2812 0.8644
1ET4 Vitamin B12 binding RNA aptamer 0.25 0.34 0.44 0.6250 1.3457 -0.0135
1F27 Biotin-binding RNA pseudoknot 0.30 0.36 0.37 0.3437 1.6203 0.4523
1L2X Viral RNA pseudoknot 0.24 0.45 0.48 0.4721 1.3181 0.5692
1LNT RNA internal loop of SRP 0.33 0.27 0.53 0.5843 1.4337 0.1379
1Q9A Sarcin/ricin domain from E.coli 23S rRNA 0.27 0.40 0.41 0.5044 1.3411 0.0597
4FE5 Guanine riboswitch aptamer 0.29 0.28 0.36 0.5300 1.3824 0.9116
1X9C All-RNA hairpin ribozyme 0.26 0.31 0.50 0.5000 1.3905 0.6630
1XPE HIV-1 B RNA dimerization initiation site 0.27 0.24 0.40 0.7037 1.2177 0.7768
2GCS Pre-cleavage state of glmS ribozyme 0.25 0.26 0.44 0.5078 1.3053 0.4062
2GDI Thiamine pyrophosphate-specific riboswitch 0.25 0.38 0.48 0.6500 1.2363 -0.0251
2OEU Junctionless hairpin ribozyme 0.23 0.30 0.37 0.9519 1.0913 0.7768
2R8S Tetrahymena ribozyme P4-P6 domain 0.27 0.36 0.53 0.5689 1.1881 0.7281
354D Loop E from E. coli 5S rRNA 0.28 0.35 0.55 0.4410 1.4938 0.0430
Overall recovery: 0.27 0.32 0.45 0.5682
Figure 12: Retrospective study of gRNAde for ranking ribozyme mutant fitness (t1 subunit). Using the backbone structure and mutational fitness landscape data from an RNA polymerase ribozyme (McRae et al., 2024), we retrospectively analyse how well we can rank variants at multiple design budgets using random selection vs. gRNAde’s perplexity for mutant sequences conditioned on the backbone structure (scaffolding subunit t1). gRNAde performs better than single site saturation mutagenesis, even when all single mutants are explored (total of 403 single mutants, 17,027 double mutants for the scaffolding subunit t1 in McRae et al. (2024)). See Section 4.3 for results on catalytic subunit 5TU and further discussions.
Appendix F Additional Figures
Figure 13: gRNAde model architecture.
@@ Appendix F Additional Figures
Figure 15: RNASolo data statistics.